### v1

# Benchmarking LLM Faithfulness in RAG with Evolving Leaderboards

Abstract
Hallucinations remain a persistent challenge for LLMs. RAG aims to reduce hallucinations by grounding responses in contexts. However, even when provided context, LLMs still frequently introduce unsupported information or contradictions. This paper presents our efforts to measure LLM hallucinations with a focus on summarization tasks, assessing how often various LLMs introduce hallucinations when summarizing documents. We discuss Vectara’s existing LLM hallucination leaderboard, based on the Hughes Hallucination Evaluation Model (HHEM). While HHEM and Vectara’s Hallucination Leaderboard have garnered great research interest, we examine challenges faced by HHEM and current hallucination detection methods by analyzing the effectiveness of these methods on existing hallucination datasets. To address these limitations, we propose FaithJudge, an LLM-as-a-judge approach guided by few-shot human hallucination annotations, which substantially improves automated LLM hallucination evaluation over current methods. We introduce an enhanced hallucination leaderboard centered on FaithJudge  [^1] https://github.com/vectara/FaithJudge, alongside our current hallucination leaderboard [^2] https://github.com/vectara/hallucination-leaderboard, enabling more reliable benchmarking of LLMs for hallucinations in RAG.
Benchmarking LLM Faithfulness in RAG with Evolving Leaderboards
Manveer Singh Tamber1∗, Forrest Sheng Bao2, Chenyu Xu2,3, Ge Luo2, Suleman Kazi2, Minseok Bae4∗, Miaoran Li3∗, Ofer Mendelevitch2, Renyi Qu2, Jimmy Lin1 1 University of Waterloo  2 Vectara  3 Iowa State University  4 Stanford University Correspondence: mtamber@uwaterloo.ca, {forrest, suleman}@vectara.com
11footnotetext: Work done while at Vectara
1 Introduction
LLMs excel in various tasks, but frequently produce hallucinations, generating false or misleading information unsupported by provided contexts or world knowledge Ji et al. (2023); Huang et al. (2025); Lin et al. (2022); Tang et al. (2023). While Retrieval-Augmented Generation (RAG) approaches Guu et al. (2020); Lewis et al. (2020b); Shuster et al. (2021) attempt to mitigate hallucinations by grounding responses in external trusted contexts, they do not fully eliminate hallucinations, as LLMs often introduce details unsupported by retrieved contexts, misrepresent information, or generate outright contradictions Niu et al. (2024).
An ongoing challenge within RAG is ensuring context-faithfulness Niu et al. (2024); Jia et al. (2023); Ming et al. (2024). Detecting when LLMs deviate from the information in the provided context remains a difficult problem. Although there has been progress, hallucination detection methods, including fine-tuned detectors largely for evaluating summaries Zhou et al. (2021); Gekhman et al. (2023); Honovich et al. (2022); Zha et al. (2023); Tang et al. (2024a) and LLM-as-a-judge techniques Zheng et al. (2023); Luo et al. (2023); Jacovi et al. (2025), continue to struggle with the accurate identification of LLM-generated hallucinations.
In this paper, we study and aim to improve hallucination evaluation in RAG by building on prior work in summary consistency evaluation. We analyze the capabilities and limitations of current hallucination detection methods, including fine-tuned models such as Vectara’s Hughes Hallucination Evaluation Model (HHEM) Bao et al. (2024) and zero-shot methods using LLM judges.
To overcome the challenges of LLM-as-a-judge techniques for zero-shot hallucination detection and fine-tuned hallucination detection models, we introduce FaithJudge, an LLM-as-a-judge approach guided by few-shot human annotations of hallucinations. FaithJudge leverages labelled hallucinations from diverse LLM generations to automate the evaluation of LLMs on their propensity to hallucinate when summarizing the same articles or using the same articles to respond to queries. This approach results in notably higher agreement with human judgments compared with existing automated methods. Additionally, we introduce an enhanced hallucination leaderboard based on FaithJudge, enabling more reliable benchmarking of hallucinations in LLM-generated summaries and responses.
We discuss both Vectara’s existing hallucination leaderboard Hughes and Bae (2023) based on HHEM and our new leaderboard based on FaithJudge. Hallucinations within RAG remain frequent and problematic, even in leading LLMs. Our approach contributes toward more accurate hallucination evaluation, aiding the development of more trustworthy generative AI systems.
Vectara serves customers across diverse industries. For many of these customers, addressing hallucinations in LLM outputs is a critical priority. Driven by this customer-centric need, we developed our hallucination leaderboard and benchmarking methods, which are presented in this paper.
2 Background
Accurate hallucination detection is essential for reliably quantifying hallucination rates in LLMs. Numerous datasets have been developed for evaluating hallucinations in summarization tasks. Earlier datasets, such as SummaC Laban et al. (2022) and AggreFact Tang et al. (2023), aggregated multiple resources, standardized labels, and classification taxonomies. However, these primarily focused on summaries from pre-ChatGPT models like fine-tuned T5 Raffel et al. (2020), BART Lewis et al. (2020a), and PEGASUS Zhang et al. (2020) models, potentially limiting their relevance to contemporary LLMs that may produce more nuanced and difficult to identify hallucinations.
Recent benchmarks address this limitation by incorporating summaries generated by modern LLMs. TofuEval Tang et al. (2024b) provided hallucination labels on topic-focused dialogue summarization tasks with LLMs including GPT-3.5-Turbo, Vicuna Chiang et al. (2023) and WizardLM Xu et al. (2023). Similarly, HaluEval Li et al. (2023) included ChatGPT-generated hallucinations across summarization, question answering (QA), and dialogue tasks, while RAGTruth Niu et al. (2024) also annotated responses from models including GPT-3.5, GPT-4 OpenAI (2023), Llama-2 Touvron et al. (2023), and Mistral Jiang et al. (2023). FaithBench Bao et al. (2025) presents human annotations of challenging hallucinations in summaries from 10 modern LLMs from 8 different model families (detailed further in Section 4).
Due to limited large-scale, human-annotated data for training hallucination detectors, early detection methods relied heavily on natural language inference (NLI) or question-answering (QA) systems Fabbri et al. (2022). For instance, SummaC aggregated sentence-level NLI entailment scores between document-summary sentence pairs. AlignScore Zha et al. (2023) extended this by training detection models on multiple semantic alignment tasks evaluated at the chunk level. MiniCheck Tang et al. (2024a) addressed data scarcity by synthesizing hallucinated examples using GPT-4 for model training.
Modern LLMs’ strong zero-shot instruction-following capabilities have also enabled LLM-as-a-judge methods Zheng et al. (2023); Luo et al. (2023); Jacovi et al. (2025); Gao et al. (2023). Instead of evaluating entire generated summaries, approaches like FACTSCORE Min et al. (2023) and RAGAS Es et al. (2024) decompose summaries into claims for granular hallucination detection.
Like Vectara’s Hallucination Leaderboard Hughes and Bae (2023), other efforts like FACTS Grounding Jacovi et al. (2025) and Galileo’s Hallucination Index Galileo (2023) also provide leaderboards to benchmark hallucinations in LLMs. Galileo’s Hallucination Index employs GPT-4o as a single LLM judge, whereas FACTS Grounding ensembles evaluations from three different LLM judges: GPT-4o, Claude-3.5-Sonnet, and Gemini-1.5-Pro.
Nonetheless, hallucination detection remains challenging, with modest effectiveness observed across current methods. Benchmarks such as AggreFact, RAGTruth, TofuEval, and FaithBench consistently show limitations in existing hallucination detectors, including LLM-based methods. Notably, FaithBench highlighted that even the best current models achieve near 50% accuracy. Both RAGTruth and TofuEval further suggest that smaller, fine-tuned detection models can perform competitively with or even outperform LLM-based evaluation approaches.
AggreFact-SOTA RAGTruth-Summ TofuEval-MB FaithBench Average
Method # Params Acc (%) F1 (%) Acc (%) F1 (%) Acc (%) F1 (%) Acc (%) F1 (%) Acc (%) F1 (%)
Claim-wise
   Fine-Tuned Hallucination Detection Models
    HHEM-1.0-Open 184M 76.0 71.0 66.2 52.2 54.4 49.9 59.3 58.8 64.0 58.0
    HHEM-2.1-Open 110M 73.2 69.7 67.7 56.1 60.9 61.2 66.7* 63.7* 67.1 62.7
    AlignScore-base 125M 69.5 61.9 60.2 42.4 51.7 44.0 60.6 59.9 60.5 52.1
    AlignScore-large 355M 73.9 69.3 67.9 54.1 56.1 52.9 62.8 59.6 65.2 59.0
    MiniCheck-Roberta-L 355M 75.7 72.5 70.5 58.6 67.6 68.5 61.6 60.0 68.8 64.9
    Bespoke-MiniCheck 7B 74.3 70.1 73.3 62.9 76.9 78.4 60.1 58.0 71.2 67.3
    TrueTeacher 11B 71.8 70.5 56.5 56.1 58.5 58.4 59.8* 51.8* 61.7 59.2
   Zero-Shot Hallucination Detection with LLMs
    RAGAS Prompt
     Qwen-2.5 7B 71.1 69.0 68.2 64.4 64.3 57.7 57.9 51.3 65.4 60.6
     Qwen-2.5 72B 74.4 69.9 75.3 64.1 69.2 70.6 64.3 57.3 70.8 65.5
     Llama-3.1 8B 69.7 65.9 68.5 59.9 72.6 74.4 60.3 57.9 67.8 64.5
     Llama-3.3 70B 77.3 74.9 80.0 75.1 73.2 70.6 58.9 49.6 72.3 67.5
     GPT-4o ? 75.9 70.3 75.7 63.7 75.2 76.7 65.3 59.0 73.0 67.4
     o3-mini-high ? 77.3 72.6 74.6 62.9 69.2 70.6 67.4 60.7 72.1 66.7
Summary-wise
   Fine-Tuned Hallucination Detection Models
    HHEM-1.0-Open 184M 78.9 79.7 53.4 51.4 56.5 39.8 50.5 40.1 59.8 52.7
    HHEM-2.1-Open 110M 76.6 76.2 64.4 67.1 69.4 62.1 52.6* 32.9* 65.8 59.6
    AlignScore-base 125M 73.8 73.9 57.6 58.2 65.6 52.8 51.3 33.8 62.1 54.7
    AlignScore-large 355M 72.7 74.2 52.8 49.6 57.4 39.2 50.3 26.1 58.3 47.3
    MiniCheck-Roberta-L 355M 74.2 72.1 66.3 60.9 54.4 45.4 55.0 53.2 62.5 57.9
    Bespoke-MiniCheck 7B 79.9 80.4 79.4 77.1 78.8 78.6 55.7 47.3 73.5 70.8
    TrueTeacher 11B 77.6 78.4 61.6 62.8 57.4 39.2 53.3* 36.7* 62.5 54.3
   Zero-Shot Hallucination Detection with LLMs
    FACTS Grounding Prompt
     Qwen-2.5 7B 66.9 68.7 61.5 63.4 62.8 54.4 52.6 33.5 60.9 55.0
     Qwen-2.5 72B 71.6 73.7 74.0 77.5 68.8 58.8 55.2 35.5 67.4 61.4
     Llama-3.1 8B 55.5 55.5 62.9 62.7 55.3 54.5 60.9 49.7 58.6 55.6
     Llama-3.3 70B 79.3 78.1 81.6 74.9 70.1 71.3 66.6 58.4 74.4 70.7
     GPT-4o ? 81.6 78.8 82.6 76.6 76.3 76.0 65.9 56.2 76.6 71.9
     o3-mini-high ? 82.1 77.8 79.8 70.6 69.2 70.6 68.8 60.7 75.0 69.9
     Luo et al. Prompt
     Qwen-2.5 7B 72.8 73.5 67.6 70.2 69.0 66.3 53.4 39.0 65.7 62.2
     Qwen-2.5 72B 78.4 78.0 81.3 81.1 83.4 80.0 58.3 44.3 75.3 70.8
     Llama-3.1 8B 60.8 51.2 63.7 52.1 57.1 55.8 51.3 51.0 58.2 52.5
     Llama-3.3 70B 79.2 79.1 81.3 82.9 73.6 66.5 58.8 43.6 73.2 68.0
     GPT-4o ? 80.4 77.5 85.1 80.9 81.6 78.7 62.5* 50.6* 77.4 71.9
     o3-mini-high ? 82.6 80.9 83.2 80.6 75.6 73.7 63.3 49.8 76.2 71.2
Table 1: Balanced Accuracy and F1-Macro of hallucination detection methods across four datasets. The final two columns report the simple average across the four datasets. We note that certain models marked with an asterisk (*) were used to select articles for the challenging FaithBench dataset.
3 Vectara’s Hallucination Leaderboard
In 2023, Vectara’s Hallucination Leaderboard Hughes and Bae (2023) was released using Vectara’s hallucination-detection model, HHEM-1.0-open. This model was later updated to HHEM-2.0 with stronger effectiveness, the ability to handle longer contexts, and multilingual capabilities. The current leaderboard relies on the open version, HHEM-2.1-open, publicly released on HuggingFace [^3] https://huggingface.co/vectara/hallucination_evaluation_model. To date, HHEM has been downloaded over 3.5 million times, reflecting strong community interest and adoption. While specific training details remain confidential, we note that HHEM-2.1-open was trained using the RAGTruth training set among other datasets.
To build Vectara’s Hallucination Leaderboard, articles were selected from diverse sources such as BBC News, CNN, Wikipedia, and the Daily Mail, following prior work on summarization evaluation and factuality verification Narayan et al. (2018); Maynez et al. (2020); Schuster et al. (2021); Thorne et al. (2018); Fabbri et al. (2021); Huang et al. (2020); Pagnoni et al. (2021); Hermann et al. (2015). Articles containing objectionable or explicit content, which LLMs may refuse to summarize, were specifically excluded. The resulting dataset comprised articles with a median length of approximately 217 words (25th percentile: 42 words; 75th percentile: 424 words).
LLMs are evaluated by prompting them to generate concise summaries strictly grounded in the provided passages. HHEM then assesses the proportion of summaries generated by the LLM containing hallucinations. Refusals are tracked by measuring the proportion of short responses (5 words or fewer). Users are also invited to submit specific models for evaluation. Continuously updated, the leaderboard now benchmarks hallucination rates of over 130 different LLMs, typically evaluating new models as soon as they become publicly available to track ongoing advances in the field.
4 FaithBench
FaithBench Bao et al. (2025) examined hallucinations in LLM-generated summaries and assessed the effectiveness of hallucination detection methods through human annotations. It includes summaries from ten state-of-the-art LLMs, including GPT-4o, GPT-3.5, Claude-3.5-Sonnet, Gemini-1.5-Flash Gemini Team (2024), and open-source models like Llama-3.1 Grattafiori et al. (2024), revealing that hallucinations remain frequent and detection methods often fail to identify them accurately.
Human annotators labelled hallucinations as Unwanted when the summary contained contradictory or unsupported information, Benign when the information was supported by world knowledge, but absent from the article, or Questionable when the classification was unclear.
Articles in FaithBench were selected from Vectara’s Hallucination Leaderboard based on frequent disagreements on summaries among hallucination detection models. True-NLI, TrueTeacher, HHEM-2.1-open, and GPT-4o/GPT-3.5 judges using the CoT prompt from Luo et al. (2023) were used to identify articles where summary hallucination classifications were most disagreed upon. The dataset includes 75 articles, each with ten annotated summaries from different LLMs.
FaithBench
Method # Params Acc (%) F1 (%)
FaithJudge Prompting
   Qwen-2.5 7B 71.9 66.6
   Qwen-2.5 72B 73.2 73.0
   Llama-3.1 8B 60.8 61.0
   Llama-3.3 70B 77.5 77.8
   GPT-4o ? 79.5 81.1
   o3-mini-high ? 84.0 82.1
   Majority Vote (Qwen 72B, Llama 70B, GPT-4o) 80.7 81.3
   Majority Vote (Qwen 72B, Llama 70B, GPT-4o, o3) 80.2 81.3
Table 2: Balanced Accuracy and F1-Macro scores for FaithJudge on FaithBench using different LLM judges. With Majority Vote, we break ties by defaulting to a classification of inconsistent.
5 FaithJudge
Human annotation is the gold standard for hallucination detection, but it is time-consuming and expensive. FaithJudge offers a scalable alternative by leveraging hallucination annotations to guide an LLM judge in evaluating new summaries. We also expand FaithJudge to other RAG tasks, including question-answering (QA) and writing overviews from structured data in the JSON format using the RAGTruth dataset Niu et al. (2024). This is detailed further in the Appendix A.
To assess a summary, FaithJudge involves prompting an LLM judge with other summaries of the same article, along with their corresponding hallucination annotations. These annotations include hallucination spans, source references, and labels of either Benign, Unwanted, or Questionable, identified by human annotators.
To evaluate the effectiveness of FaithJudge, we use the fact that each FaithBench article has summaries from ten different LLMs. The judge is given the other nine annotated summaries as context, and its assessments on each summary from FaithBench are compared to human annotations. As shown in Section 6, FaithJudge substantially improves automated hallucination evaluation, outperforming existing detection methods by leveraging human-labelled examples. This allows for more accurate automated hallucination evaluation, where existing hallucination detection methods continue to lag.
6 Evaluating Hallucination Detectors
6.1 Evaluation Datasets
We evaluate leading hallucination detection methods on four datasets: FaithBench, AggreFact Tang et al. (2023), RAGTruth Niu et al. (2024), and TofuEval-MeetingBank Tang et al. (2024b). While each of these datasets has previously analyzed hallucination detection individually, we provide a broader comparison across all four, motivating the need for our FaithJudge approach.
For FaithBench, we assign each summary the most severe hallucination label given by a majority of the annotators. We evaluate using summaries labelled either Unwanted or Consistent, excluding Benign and Questionable cases due to their more ambiguous nature. This slightly differs from the original FaithBench evaluation, which pooled the worst label across all annotators for each summary and combined Benign cases with Consistent ones, while combining Unwanted cases with Questionable ones for the binary classification problem.
For AggreFact, we evaluate on the SOTA subset of summaries, which involves annotated summaries generated by fine-tuned T5 Raffel et al. (2020), BART Lewis et al. (2020a), and PEGASUS Zhang et al. (2020) models. For RAGTruth, we evaluate only on the annotated summaries subset. Lastly, for TofuEval-Meetingbank, we evaluate on summaries generated using articles from the MeetingBank dataset Hu et al. (2023).
Figure 1: Proportion of summary FaithBench labels (left) and FaithJudge predictions (right) across models. For FaithBench labels, red indicates Unwanted, orange indicates Questionable, yellow indicates Benign, while green indicates consistent. For FaithJudge predictions, red indicates hallucinated, and green indicates consistent summaries. Each bar shows the proportion of summaries falling into each category.
6.2 Existing Hallucination Detectors
Table 1 compares the effectiveness of fine-tuned hallucination detectors and zero-shot LLM-based methods across various datasets. We evaluate Vectara’s HHEM models alongside AlignScore, MiniCheck, including Bespoke-MiniCheck Bespoke (2024), and TrueTeacher. We also include current LLMs, such as GPT-4o and o3-mini (high reasoning), as well as open-source models Qwen2.5 (7B and 72B), Llama-3.1 (8B), and Llama-3.3 (70B). The o3-mini model, in particular, excels in reasoning tasks.
Classification methods are separated into claim-wise and summary-wise classification. Claim-wise evaluation involves decomposing sentences from summaries into individual claims using Llama-3.3 (70B) and a similar prompt from Tang et al. (2024a), while summary-wise methods assess the entire summary at once.
For LLM-based detection, we test three prompts: (1) the RAGAS prompt, which verifies lists of claims, (2) the FACTS Grounding JSON prompt, shown to be the most effective of the prompts tested in  Jacovi et al. (2025) for GPT-4o, and (3) the CoT-based prompt from Luo et al. (2023). We modify prompts slightly as needed for clearer final outputs and to specifically evaluate summaries.
Binary Classification
Gold Truth Consistent Inconsistent
Unwanted $7474$ $322322$
Questionable $2929$ $3838$
Benign $5050$ $3434$
Consistent $176176$ $2727$
Ternary Classification
Gold Truth Consistent Benign Unwanted
Unwanted $8484$ $1818$ $294294$
Questionable $2828$ $1313$ $2626$
Benign $5151$ $1010$ $2323$
Consistent $179179$ $44$ $2020$
Table 3: Confusion matrices for FaithJudge prompted for classification on FaithBench summaries.
Table 1 shows that, similar to findings in previous work, hallucination detection remains challenging. We note that certain models marked with an asterisk (*) were used to select articles for the adversarially challenging FaithBench dataset. Consequently, these models, including HHEM, may perform worse on FaithBench in summary-wise classification than they otherwise would. Zero-shot classification using GPT-4o and o3-mini-high tends to perform best, both using summary-wise classification with either the FACTS Grounding JSON prompt or the Luo et al. (2023) prompt. However, their average effectiveness remains modest, with balanced accuracy below 78% and F1-macro below 72%. Considering FaithBench, the highest balanced accuracy is achieved by o3-mini-high at 68.8% while the highest F1-macro of 63.7% is achieved by the HHEM model when considering claim-wise classification.
The table illustrates improved effectiveness with increased model size: larger open-source models generally outperform smaller ones, and GPT-4o and o3-mini-high achieve the highest overall effectiveness. However, although HHEM-2.1-open is the smallest model tested, it performs strongly, outperforming several larger models. Among the fine-tuned models, only the 7B-parameter MiniCheck achieves higher average scores for summary-wise classification, while both MiniCheck variants outperform it in claim-wise classification.
Overall, fine-tuned models can score stronger than smaller prompted LLMs, but the largest LLMs typically yield the best results, even while being zero-shot methods. Regardless, the examined methods demonstrate modest effectiveness in general, with particularly weak effectiveness on FaithBench, which captures a diverse set of LLM summaries but is designed to be challenging for hallucination detection models.
Figure 2: Sensitivity and specificity with FaithJudge as the number of examples in the prompt are increased. We place an asterisk (*) next to the 10 because, in this case, FaithJudge is shown annotations for the summary it is evaluating.
Figure 3: Comparison of LLM rankings across FaithBench (based on (U) Unwanted, (B) Benign, and (Q) Questionable hallucination annotations), FaithJudge, and Vectara’s Hallucination Leaderboard. Rankings reflect the number of hallucinated summaries (from least to most).
6.3 FaithJudge
Table 2 presents the effectiveness of FaithJudge on FaithBench using various LLMs. The highest effectiveness is achieved using the o3-mini-high judge, reaching a balanced accuracy of 84% and an F1-macro of 82.1%, allowing for higher agreement with human annotation on FaithBench than the existing methods discussed in Table 1. Although the effectiveness of FaithJudge is not perfect, this may be partly explained by disagreements in human annotation. While human annotation is the gold standard, the FaithBench paper Bao et al. (2025) noted imperfect inter-annotator agreement in general and low inter-annotator agreement on more gray-area Benign and Questionable Hallucinations. We also note that we observe some erroneous annotations within FaithBench.
Effectiveness generally improves with increasing model size. We also tested an ensemble approach inspired by FACTS Grounding but found that combining predictions from multiple models, including o3-mini-high itself, did not outperform o3-mini-high alone. Therefore, we adopt the o3-mini-high judge as the standard for FaithJudge, with the possibility of using a stronger LLM judge down the line.
Figure 1 displays the distribution of FaithJudge predictions across LLMs. While effective, FaithJudge with o3-mini-high tends to underpredict hallucinations. This is evident for Command-R, Mistral, and Qwen, where fewer summaries were flagged as hallucinated compared to the number labelled Unwanted by annotators in FaithBench.
Table 3 presents confusion matrices for both binary and ternary classification using FaithJudge. We observe that Benign summaries, in particular, are difficult for FaithJudge to classify correctly. In the ternary setting, FaithJudge often misclassifies Benign summaries, generally labelling them as Consistent. Similarly, Questionable summaries are classified unreliably, though this aligns with expectations. For this reason, we only employ FaithJudge for binary classification.
Finally, Figure 2 shows the sensitivity and specificity of FaithJudge as the number of annotated examples provided increases. Specificity remains consistently high, though slightly decreasing as more examples are given, while sensitivity notably improves as the number of examples increases. This indicates that providing more annotated examples causes FaithJudge to predict hallucinated cases more often and better identify hallucinations.
7 Leaderboard Rankings
Figure 3 compares the ranking of the 10 LLMs studied in FaithBench based on human-annotated hallucinations with rankings from FaithJudge and Vectara’s existing hallucination leaderboard.
The left-most plot shows that rankings vary depending on the type of hallucination considered: Unwanted, Benign, or Questionable, even when assessed by human annotation. The other plots show that when considering all types of hallucination annotations, rankings in FaithBench align more closely with FaithJudge than with the existing leaderboard. FaithJudge rankings show six inversions compared to rankings from FaithBench considering Unwanted, Benign, and Questionable hallucinations, while the existing leaderboard rankings using HHEM shows 16 inversions.
8 Conclusion
In this paper, we presented our efforts at Vectara in evaluating and benchmarking hallucinations in RAG, discussing and building on our established hallucination leaderboard, and proposing FaithJudge. We identified effectiveness limitations in existing hallucination detection methods, including our own HHEM model. To address these challenges, we proposed FaithJudge, an approach that leverages human hallucination annotations to enhance automated hallucination detection, achieving greater effectiveness, but requiring annotations from summaries of the same articles.
Beyond FaithBench, we extend FaithJudge to additional RAG tasks, including question answering and data-to-text generation, using annotated examples of hallucinations from the RAGTruth dataset. We discuss this further in Appendix A. We also apply FaithJudge to a broader set of LLMs, producing leaderboard-style rankings that currently include 30 models. We share some of these results in Appendix C, providing a framework for more accurate faithfulness evaluation across diverse models and RAG tasks. We hope to continue to update our leaderboard to evaluate new models and to use improved LLM judges.
Acknowledgements
We respectfully acknowledge the late Simon Mark Hughes, who led the development of the original HHEM model and Vectara’s Hallucination Leaderboard. His contributions laid important groundwork for Vectara’s ongoing research and continue to leave a lasting influence on our work.
Limitations
There are some limitations with our evaluation methodology. First, our evaluation focuses exclusively on faithfulness and does not address the overall quality or usefulness of summaries and answers. Though summary and answer quality are important in RAG applications, we consider this evaluation somewhat orthogonal to faithfulness.
One issue to consider is that an extractive summarizer or an LLM that simply copies parts of or the entire article in its response would technically avoid hallucinations. Nonetheless, we maintain that evaluating LLMs through hallucinations in generated summaries is promising because these hallucinations remain persistent.
Finally, while the o3-mini-high judge demonstrates strong effectiveness, there remains room for enhancing accuracy and agreement with human annotators. We hope that as LLMs continue to improve, replacing o3-mini-high in FaithJudge would allow for more accurate and reliable evaluation.
Appendix A Adding More Evaluation Tasks
Judged by o3-mini-high Judged by gemini-2.0-flash Judged by llama-4-maverick
Evaluated Model Hallucinated Responses Rank Hallucinated Responses Rank Hallucinated Responses Rank
gemini-2.0-flash 52 1 31 2 71 1
o3-mini-high 64 2 29 1 94 2
llama-4-maverick 105 3 72 3 110 3
Table 6: Total number of hallucinated responses per evaluated model, as judged by each model. Rankings indicate relative effectiveness in terms of hallucination frequency, from least to most.
Rank Model Overall Hallucination Rate FaithBench (Summary) RAGTruth (Summary) RAGTruth (QA) RAGTruth (Data-to-Text)
1 gemini-2.5-pro-exp 7.63% 18/72 14/150 1/139 6/150
2 gemini-2.0-flash 10.18% 21/72 10/150 1/139 20/150
3 gpt-4.5-preview 11.94% 27/72 15/150 7/139 12/150
4 o3-mini-high 12.52% 25/72 12/150 9/139 18/150
5 gpt-3.5-turbo 14.87% 32/72 13/150 8/139 23/150
6 gpt-4o 15.85% 29/72 15/150 7/139 30/150
7 claude-3.7-sonnet 16.05% 28/72 22/150 13/139 19/150
8 llama-3.3-70b 16.44% 32/72 13/150 6/139 33/150
9 phi-4 17.03% 32/72 12/150 6/139 37/150
10 mistral-small-24b 17.03% 31/72 15/150 14/139 27/150
11 llama-4-maverick 20.55% 37/72 20/150 13/139 35/150
12 llama-3.1-8b 28.38% 32/72 19/150 17/139 77/150
Table 7: FaithJudge rankings for 12 LLMs, based on the number of hallucinated responses across four evaluation subsets: article summarization from FaithBench and RAGTruth, as well as question answering and data-to-text writing from RAGTruth.
While FaithBench provides hallucination annotations across 10 different LLMs, it is limited to evaluating summaries only.
To broaden the scope of FaithJudge beyond summarization, we incorporate annotated responses from the RAGTruth dataset Niu et al. (2024). RAGTruth includes three types of tasks: Summarization, Question Answering (QA), and a Data-to-Text generation task that requires generating an overview of a business from JSON data sourced from the Yelp Open Dataset. It contains human-annotated hallucination labels for responses generated by six LLMs: GPT-3.5, GPT-4 OpenAI (2023), Llama-2 (7B, 13B, 70B) Touvron et al. (2023), and Mistral-7B Jiang et al. (2023).
For each RAGTruth task, we take up to 150 sources (articles for summarization, queries and passages for question-answering, and JSON data for data-to-text) with their corresponding annotated responses from the test set first and then the dev set. We remove sources where none of the LLM responses have a hallucination annotation.
Table 4 compares the effectiveness of FaithJudge against the zero-shot FACTS Grounding JSON prompt previously shown as an effective prompt in Jacovi et al. (2025), on the FaithBench and RAGTruth subsets used in our leaderboard. In each setting, FaithJudge achieves stronger agreement with human hallucination annotations, highlighting its strength across tasks beyond summarization.
Appendix B Judge Bias
The FACTS Grounding leaderboard Jacovi et al. (2025) uses three different LLM judges to mitigate bias arising from any single judge favoring its own outputs. Inspired by this, we analyze judge bias using Tables 5 and 6, which evaluate the impact of using different judges across all subsets included in our leaderboard.
Table 5 reports the effectiveness of three different LLMs when used as judges. The table shows that o3-mini-high remains a relatively effective LLM for FaithJudge, often scoring the highest. The table also shows that using multiple judges can improve effectiveness further, though we note that in some cases, individual LLMs can score higher than the majority vote approach between the three LLMs. For example, o3-mini-high scores higher than the ensembling approach when evaluating on the RAGTruth QA subset.
Table 6 explores how each judge model ranks other LLMs. Interestingly, o3-mini-high and llama-4-maverick both rank gemini-2.0-flash as having the fewest hallucinated responses, while gemini-2.0-flash ranks itself second to o3-mini-high, with only a small difference in counts (29 vs. 31).
While using multiple judges might enhance robustness and reduce individual model bias, we currently rely on a single judge to reduce computational costs. As stronger LLMs become available, we plan to update FaithJudge by substituting the current judge model with a more effective one.
Appendix C Leaderboard Rankings
Table 7 presents FaithJudge rankings for a range of LLMs. In addition to detecting hallucinations, we also prompt FaithJudge to flag responses that are invalid, for example, when a model fails to meaningfully summarize an article. For simplicity, we count these as hallucinated responses. Models are ranked based on their overall hallucination rate, calculated as the total number of hallucinated or invalid responses across all four evaluation subsets. We plan to continue evaluating LLMs using FaithJudge alongside the existing leaderboard.